John slides down a hill on a toboggan. He starts from rest and speeds up to 7.9 m/s in 8.0 s.
Find his acceleration in m/s2
.
a. a. 0.66 m/s2
[downhill]
b. b. 0.77 m/s2
[downhill]
c. c. 0.88 m/s2
[downhill]
d. d. 0.99 m/s2
[downhill]
1
Expert's answer
2020-05-21T09:32:07-0400
"a=\frac{v}{t}=\frac{7.9}{8}=0.99\frac{m}{s^2}"

Answer: d. 0.99 m/s2

[downhill]